# FinAudio: A Benchmark for Audio Large Language Models in Financial Applications

### v1

@@ Abstract
Audio Large Language Models (AudioLLMs) have received widespread attention and have significantly improved performance on audio tasks such as conversation, audio understanding, and automatic speech recognition (ASR). Despite these advancements, there is an absence of a benchmark for assessing AudioLLMs in financial scenarios, where audio data, such as earnings conference calls and CEO speeches, are crucial resources for financial analysis and investment decisions. In this paper, we introduce FinAudio, the first benchmark designed to evaluate the capacity of AudioLLMs in the financial domain. We first define three tasks based on the unique characteristics of the financial domain: 1) ASR for short financial audio, 2) ASR for long financial audio, and 3) summarization of long financial audio. Then, we curate two short and two long audio datasets, respectively, and develop a novel dataset for financial audio summarization, comprising the FinAudio benchmark. Then, we evaluate seven prevalent AudioLLMs on FinAudio. Our evaluation reveals the limitations of existing AudioLLMs in the financial domain and offers insights for improving AudioLLMs. All datasets and codes will be released.
@@ 1 Introduction
Audio data plays a critical role in the financial domain, encompassing earnings conference calls, investor presentations, and customer service interactions (see Figure 1). Timely and accurate interpretation of financial audio data underpins financial services, sentiment analysis, and investment decisions Qin and Yang (2019); Yang et al. (2020); Cao et al. (2024).
Recent developments in financial AI have been propelled by breakthroughs in natural language processing (NLP), particularly large language models (LLMs), such as BloombergGPT Wu et al. (2023b), PIXIU Xie et al. (2023), FinGPT Liu et al. . Financial evaluation suites like FinBen Xie et al. (2025) further provide extensive assessments of LLMs in financial NLP tasks. Despite these advancements, there remains a critical gap: there is no audio-focused financial LLM or a corresponding evaluation benchmark. Although multimodal financial LLMs (e.g., FinLLaVA-8B Xie et al. (2024)) can process text, images, and charts, they cannot yet handle audio data.
Figure 2: Overview of FinAudio benchmark.
In the general domain, AudioLLMs have made significant progress Wu et al. (2023a); Zhang et al. (2024). AudioLLMs extend foundational LLMs with an audio encoder that aligns audio inputs to the LLM’s input space Radford et al. (2023); Hu et al. (2024); Chu et al. (2023, 2024); Tang et al. , enabling audio-centric tasks such as automatic speech recognition (ASR), speech question answering (SQA), audio scene understanding, and voice analysis. While their effectiveness in these tasks has been explored in a variety of general scenarios Yang et al. (2024); Wang et al. (2024), a comprehensive benchmark for financial audio remains absent. This gap limits the research community’s ability to systematically evaluate performance, compare strategies, and identify biases in financial audio datasets, hindering downstream applications such as investment advisory and financial services.
@@
To address this critical need, we introduce FinAudio, the first AudioLLM benchmark tailored for the financial domain. Specifically, FinAudio includes three tasks representating real-world financial scenarios: (1) ASR for short financial audio clips, (2) ASR for long financial audio recordings, and (3) summarization of financial audio data. For this purpose, we have collected four open-source financial audio datasets. Two of these datasets comprise numerous short audio clips (each under one minute), while the other two feature much longer recordings, about 45 to 60 minutes, which is sufficient to include a complete earnings conference call. Additionally, We collect a new dataset tailored for financial audio summarization, named FinAudioSum. Altogether, these five datasets form the FinAudio benchmark, comprising over 400 hours of financial audio. FinAudio’s scale is comparable to general-domain AudioLLM benchmarks.
On FinAudio, we evaluate seven representative AudioLLMs and observed the following: (a) AudioLLMs perform better in ASR tasks on short financial audio clips compared to longer recordings; (b) the performance of long-audio ASR limits audio summarization quality; (c) AudioLLMs exhibit significant variability in their robustness to different instructions; (d) The open-source AudioLLM Whisper-v3 outperforms closed-source models across all ASR datasets, offering a potential low-cost and privacy-preserving solution for both the financial industry and individual investors. Our main contributions are below:
(1) We propose FinAudio, the first comprehensive open-source evaluation benchmark for AudioLLMs in the financial audio data.
(2) We design three tasks that reflect realistic financial speech scenarios and compile five evaluation datasets, including a newly developed dataset for financial audio summarization.
(3) We conduct a systematic evaluation of seven AudioLLMs, showcasing their advantages and limitations and highlighting challenges and future research directions.
2 Related Work
2.1 Advancements in AudioLLM
Among the modalities addressed by multimodal LLMs, audio stands out as particularly critical and challenging Su et al. (2025). Several multimodal models, such as Gemini Pro Team et al. (2023) and PandaGPT Su et al. (2023), have been adapted to process audio data. Prominent AudioLLMs include AudioGPT Huang et al. (2024), SpeechGPT Zhang et al. (2023), AudioPaLM Rubenstein et al. (2023), Qwen-Audio Chu et al. (2023), WavLLM Hu et al. (2024), Whisper Radford et al. (2023), Salmonn Tang et al. , and Qwen2-Audio Chu et al. (2024). Correspondingly, general-domain benchmarks like AirBench Yang et al. (2024), emphasizing conversational scenarios, and AudioBench Wang et al. (2024), evaluating eight tasks across twenty datasets, have been developed to assess AudioLLMs. However, these benchmarks contain limited financial audio data, inadequate for a comprehensive evaluation in financial contexts.
Figure 3: Evaluation pipelines for the three FinAudio’s major tasks.
2.2 Progress in Multimodal Financial LLMs And Benchmark
Recent advances in general-domain LLMs have accelerated the emergence of specialized financial LLMs, such as FinGPT Liu et al. , InvestLM Yang et al. (2023), and BloombergGPT Wu et al. (2023b). The progression toward multimodal financial LLMs—including Open-FinLLMs Xie et al. (2024) and FinTral Bhatia et al. (2024)—further highlights efforts to address complex financial tasks involving charts and tabular data. These multimodal models overcome the limitations of traditional approaches Ang and Lim (2022); Bhatia et al. (2024); Yu and Suchow (2022) by more flexibly adapting to financial contexts. As financial LLMs become increasingly specialized, the demand for robust evaluation grows. In response, various benchmarks have been proposed. FLUE Shah et al. (2022) introduced the first benchmark for financial NLP evaluation, covering tasks such as sentiment analysis, headline classification, named entity recognition, structure boundary detection, and question answering. PIXIU Xie et al. (2023) extended these evaluations into multimodal document analysis, while FinBen Xie et al. (2025) significantly expanded coverage, featuring 36 datasets across 24 financial tasks. Despite these advances, existing benchmarks remain predominantly text-focused, overlooking audio—an essential modality in financial applications.
@@ 3 The FinAudio Benchmark
FinAudio aims to evaluate the performance of AudioLLMs on financial audio data. Our evaluation primarily focuses on the models’ ability to comprehend financial audio across various financial scenarios. Accordingly, our benchmarks are designed to evaluate three crucial aspects: 1) ASR for short financial audio clips, which are common in financial services and conversations; 2) ASR for long financial audio recordings, typically in major financial meetings; 3) Summarization for long financial audio recordings, aimed at testing the capacity of AudioLLMs to construct frameworks for understanding extensive financial audio content. For each task, we both compile existing datasets and develop new ones to ensure a comprehensive understanding ability. Figure 1 is an overview of FinAudio . This section will detail the task definitions and the development of the evaluation datasets.
@@ 3.1 FinAudio Tasks
Here we specify the FinAudio tasks, which are illustrated in Figure 3.
@@
3.1.1 ASR for short financial audio clip
Automatic Speech Recognition converts spoken language into text. Short audio clips from financial chatbot assistants or financial news often contain dense financial terms. This task evaluates AudioLLMs’ accuracy in transcribing financial audio clips, directly impacting applications such as financial voice assistants.
In this task, each input is represented as $(A,Q,R)$, where $A$ is an audio clip, $Q$ is the prompt instruction, and $R$ is the reference transcript. The transcribed text $T$ is generated as $T=\textit{AudioLLM}(A,Q)$. The Word Error Rate (WER) is computed as $\text{WER}=\frac{S+D+I}{N}$, where $S$, $D$, and $I$ represent the number of substitutions, deletions, and insertions, respectively, and $N$ is the total number of words in the reference ($N=S+D+C$, with $C$ indicating correct words). A lower WER indicates better ASR performance.
3.1.2 ASR for long financial audio recordings
Long audio recordings are common in finance, capturing events like quarterly earnings calls, mergers and acquisitions discussions, and central bank meetings. In this scenario, ASR task is thus used to transcribe lengthy financial audio, evaluating AudioLLMs’ capability to handle extended audio inputs. Challenges include long durations and complex financial terminology. This task assesses the models’ ability to consistently maintain transcription accuracy over long recordings, which is essential for comprehensive financial analysis.
To handle the varying window lengths and the inability of current AudioLLMs to process long audio files within a single window, we standardize audio inputs to 30-second segments, dividing each raw audio $A$ into chunks $\{a_{1},a_{2},\dots,a_{n}\}$. We transcribe each chunk $a_{i}$ using AudioLLMs into $t_{i}$, and concatenated the resulting transcriptions as $T=[t_{1};t_{2};\dots;t_{n}]$. The transcription quality is then evaluated by computing the WER.
3.1.3 Financial audio summarization
This task measures the ability of AudioLLMs to accurately summarize long financial audio content, a scenario unique to financial applications. It evaluates models’ understanding of financial discussions, extraction of key points, and effectiveness in condensing lengthy audio into concise summaries. This capability is crucial for stakeholders needing quick insights from extensive financial dialogues, such as investor briefings or regulatory meetings.
AudioLLMs can process audio with a maximum length of 30 seconds. Therefore, AudioLLM cannot perform the summarization task independently, we developed a processing pipeline. Given input data represented as $(A,Q,L)$, where $A$ denotes the long audio recording, $Q$ the prompt instruction, and $L$ the reference summary, the pipeline first segments the long audio into smaller chunks (see Section 3.1.2). Each chunk is transcribed by an AudioLLM into text segments $t_{i}$, which are combined into a complete transcription $T$. This transcription is then input into an LLM to generate the summary $S$. Summarization performance is evaluated using two widely adopted summarization evaluation metrics: Rouge-L Lin (2004) and BertScore Zhang et al. , computed between the generated summary ($S$) and the reference summary ($L$).
Datasets Models Whisper Qwen2-Audio Qwen2-Audio SALMONN SALMONN Gemini Gemini -v3 -7B -7B-Instruct -7B -13B -1.5-flash -2.0-flash MDRM-test 2.14 3.97 4.68 51.52 49.17 4.850 4.321 SPGISpeech-test 2.88 4.42 5.74 39.51 41.17 5.802 5.143 Earning-21 11.85 26.06 29.58 83.20 80.54 18.58 19.17 Earning-22 15.93 42.76 33.65 88.50 86.23 27.13 28.12
Table 2: Comparison of model performance based on WER across various speech datasets on ASR task.
3.2 FinAudio Datasets
The evaluation dataset in FinAudio is developed from two primary data sources: 1) existing open-source financial audio data originally created for non-LLM evaluation purposes. 2) novel datasets introduced in this work. Table 1 summarizes the key statistics. This section outlines their key information, leaving the details to Appendix A.
3.2.1 Short financial audio clip datasets
Two datasets provide financial audio clips: MDRM Qin and Yang (2019) and SPGISpeech O’Neill et al. (2021), both derived from earnings conference call recordings. The original audio data was segmented at the sentence level and aligned with corresponding transcripts. Each dataset is initially split into training and test subsets; we utilize only the test sets for evaluation. The MDRM-test set includes 22,208 audio clips totaling 87 hours, while the SPGISpeech-test set contains 39,341 clips totaling 130 hours.
3.2.2 Long financial audio recording datasets
Two datasets provide long financial audio recordings: Earnings-21 Del Rio et al. (2021), containing 44 earnings calls (approximately 39 hours), and Earnings-22 Del Rio et al. (2022), with 125 earnings calls (120 hours). Each recording includes a corresponding manual transcript. We use both datasets entirely for evaluation.
3.2.3 FinAudioSum
We create FinAudioSum based on the ECTSum dataset Mukherjee et al. (2022), originally designed for earnings call summarization using textual data. ECTSum comprises 2,425 earnings transcripts paired with expert-generated, telegram-style summaries. We obtain corresponding audio recordings for the ECTSum test set from earningcast [^1] https://earningscast.com/. Overlapping recordings with Earnings-21 and Earnings-22 (spanning 2019–2022) are removed. The final FinAudioSum dataset includes 64 recordings totaling 55 hours.
4 Model and Experiment Setup
We evaluate the performance of seven representative AudioLLMs on the FinAudio benchmark: five open-source models — Whisper Radford et al. (2023), Qwen2-Audio-7B, Qwen2-Audio-7B-instruct Chu et al. (2024), SALMONN-7B, SALMONN-13B Tang et al. ; two closed-source models — Gemini-1.5-flash and Gemini-2.0-flash Team et al. (2023). These models span various parameter sizes, open- and closed-source implementations, and include both base and instruction-tuned versions, enabling a comprehensive assessment of AudioLLMs for financial audio tasks. To ensure fairness, we standardized the prompt as: “Convert the audio speech into a text transcript,”. For summarization, transcripts from each model were summarized using GPT-4o (temperature=0). For open-sourced AudioLLMs, these experiments are exclusively conducted on two NVIDIA RTX A6000 GPUs, each equipped with 48GB of DRAM. Each round of evaluation takes approximately 200 hours to complete. For experiments in Tasks 1 and 2, we repeated evaluations three times and report the average results.
5 Results and Analysis
In this section, we first analyze the performance of seven models on two ASR tasks, then evaluate their summarization results, and finally examine their robustness to varying prompts.
@@ 5.1 Results on ASR Tasks
Table 2 shows that Whisper-v3 achieves notably low WER scores (2%–3%) on both datasets, highlighting its robust speech recognition without domain-specific tuning. The Qwen2-Audio series (base and instruct versions) and Gemini exhibit moderate performance with WER scores between 4% and 6%. Conversely, SALMONN (7B and 13B) demonstrates substantially lower effectiveness, with WER around or exceeding 40%–50%. Despite SALMONN’s limited performance, these findings highlight the potential of other mainstream AudioLLMs to support voice-based financial services, particularly conversational assistants handling iterative short audio communications.
@@
In the long financial audio ASR task, Whisper-v3 remained the top performance. It achieves lower WER in the range of approximately 12% - 16%. The Gemini series emerged as the second-best performer, with WER spanning roughly 18–28%. The error rate of Qwen2-Audio series models has also increased significantly. SALMONN once again reported the highest error rates across both financial datasets (80–88%). These results reveal a decline in performance across all AudioLLMs, indicating that AudioLLMs face significant challenges when processing long financial audio data.
The SALMONN series models performed poorly on both ASR tasks. Upon reviewing the experimental logs, we observed that SALMONN frequently produced empty outputs or failed to understand prompt instructions. This suggests that SALMONN might be overfitting to specific audio data or tasks, lacks robust instruction-following capabilities, and exhibits a limited understanding of financial audio content.
@@
In Figure 4, we find no significant difference in performance between Gemini-1.5 and Gemini-2.0. The open-source model Whisper-v3 achieves the best overall performance across all datasets. These results demonstrate robust capabilities even within long financial audio recognition. The Qwen-2 series, another open-source model, delivers comparable results to the closed-source Gemini models on short-audio ASR tasks but performs slightly worse on long-audio ASR. These findings demonstrate that open-source models can surpass closed-source counterparts in financial audio scenarios. This suggests that open-sourced AudioLLMs provide an effective, low-cost, and privacy-preserving solution for the financial industry and individual investors.
5.2 Results on summarization task

### v2

@@ Abstract
Audio Large Language Models (AudioLLMs) have received widespread attention and have significantly improved performance on audio tasks such as conversation, audio understanding, and automatic speech recognition (ASR). Despite these advancements, there is an absence of a benchmark for assessing AudioLLMs in financial scenarios, where audio data, such as earnings conference calls and CEO speeches, are crucial resources for financial analysis and investment decisions. In this paper, we introduce FinAudio, the first benchmark designed to evaluate the capacity of AudioLLMs in the financial domain. We first define three tasks based on the unique characteristics of the financial domain: 1) ASR for short financial audio, 2) ASR for long financial audio, and 3) summarization of long financial audio. Then, we curate two short and two long audio datasets, respectively, and develop a novel dataset for financial audio summarization, comprising the FinAudio benchmark. We evaluate seven prevalent AudioLLMs on FinAudio. Our evaluation reveals the limitations of existing AudioLLMs in the financial domain and offers insights for improving AudioLLMs (Project Page).
@@ 1 Introduction
Timely and accurate interpretation of financial audio underpins sentiment analysis and investment decision-making [qin2019you, yang2020html, cao2024ecc]. Financial audio data serves as a critical foundation across diverse scenarios, including earnings conference calls, investor presentations, and customer service interactions. Recent developments in financial AI have been propelled by breakthroughs in language-centric models, particularly large language models (e.g., BloombergGPT [wu2023bloomberggpt], PIXIU [xie2023pixiu], FinGPT [liu2023fingpt]) and by evaluation suites such as FinBen [xie2025finben] that benchmark LLM performance on financial NLP tasks. However, a critical gap remains: none of these models or benchmarks addresses financial audio, and even the latest multimodal model (FinLLaVA-8B [xie2024open], designed to handle text, images, and charts) cannot process audio inputs.
In the general domain, AudioLLMs incorporate an audio encoder into language models [wu2023multimodal, zhang2024mm, radford2023robust, hu2024wavllm, chu2023qwen, chu2024qwen2, tangsalmonn]. This integration allows LLMs to accept audio inputs and perform tasks such as automatic speech recognition (ASR) and spoken question answering (SQA). These audio-capable LLMs have demonstrated effectiveness on general benchmarks [yang2024air, wang2024audiobench]. Yet, the financial domain still lacks a dedicated audio benchmark, leaving researchers unable to systematically evaluate model performance on financial audio or compare different approaches, thereby hindering progress in audio-driven financial applications, such as investment advisory services.
@@
To address this critical need, we introduce FinAudio, the first AudioLLM benchmark tailored for the financial domain. Specifically, FinAudio includes three tasks representing real-world financial scenarios: (1) ASR for short financial audio clips (each under 1 minute), (2) ASR for long financial audio recordings (45–60 minutes, e.g. earnings calls), and (3) summarization of financial audio data. To support these tasks, we curated five financial audio datasets. Four are from existing open-source datasets (two with short clips and two with full-length recordings). Additionally, we created a new dataset tailored for financial audio summarization, named FinAudioSum. Altogether, the FinAudio benchmark provides 400+ hours of financial audio data, comparable in scale to general-domain AudioLLM benchmarks.
On FinAudio, we evaluate seven representative AudioLLMs and observed the following: (a) AudioLLMs achieve higher accuracy on short clips than on long recordings; (b) Errors in long-audio ASR directly undermine the quality of audio summaries. (c) AudioLLMs exhibit significant variability in their robustness to different instructions; (d) The open-source model Whisper-v3 outperforms closed-source models on all ASR tasks, offering a low-cost, privacy-preserving solution for financial applications. Our main contributions are below: (1) We present FinAudio, the first comprehensive open-source benchmark for evaluating AudioLLMs on financial audio tasks. (2) We design three realistic financial speech tasks and compile five corresponding datasets, including the new FinAudioSum dataset for audio summarization. (3) We conduct an extensive evaluation of seven AudioLLMs on these tasks, highlighting each model’s strengths and limitations and outlining challenges for future research. FinAudio aims to accelerate the development of robust AudioLLMs for financial applications, promoting transparency and accessibility of financial information processing.
@@ 2 The FinAudio Benchmark
Figure 1 is an overview of FinAudio. This section defines three key tasks in FinAudio and describes how we constructed the evaluation datasets. Each task corresponds to a specific real-world financial audio scenario, allowing us to test different aspects of an AudioLLM’s capabilities. For each task, we both compile existing datasets and develop new ones to ensure a comprehensive evaluation.
Figure 1: Evaluation pipelines for the three FinAudio’s tasks.
@@ 2.1 FinAudio Tasks
2.1.1 ASR for short financial audio clip
Automatic Speech Recognition converts spoken language into text. Short audio clips from financial chatbot assistants or financial news often contain dense financial terms. This task evaluates AudioLLMs’ accuracy in transcribing financial audio clips, directly impacting applications such as financial voice assistants. In this task, each input is represented as $(A,Q,R)$, where AA is an audio clip, $Q$ is the prompt instruction, and $R$ is the reference transcript. The transcribed text $T$ is generated as $T=\textit{AudioLLM}(A,Q)$. The Word Error Rate (WER) is computed as: $\text{WER}=\frac{S+D+I}{N}$, where $S$, $D$, and II represent the number of substitutions, deletions, and insertions, respectively, and NN is the total number of words in the reference ($N=S+D+C$, with $C$ indicating correct words). A lower WER indicates better ASR performance.
2.1.2 ASR for long financial audio recordings
Long audio recordings are common in finance, capturing events like quarterly earnings calls, mergers and acquisitions discussions, and central bank meetings. In this scenario, ASR task is thus used to transcribe lengthy financial audio, evaluating AudioLLMs’ capability to handle extended audio inputs. Challenges include long durations and complex financial terminology. This task assesses the models’ ability to consistently maintain transcription accuracy over long recordings, which is essential for comprehensive financial analysis. To handle the varying window lengths and the inability of current AudioLLMs to process long audio files within a single window, we standardize audio inputs to 30-second segments, dividing each raw audio AA into chunks $\{a_{1},a_{2},\dots,a_{n}\}$. We transcribe each chunk $a_{i}$ using AudioLLMs into $t_{i}$, and concatenated the resulting transcriptions as $T=[t_{1};t_{2};\dots;t_{n}]$. WER then evaluates the transcription quality.
2.1.3 Financial audio summarization
This task assesses the ability of AudioLLMs to accurately summarize lengthy financial audio content, a scenario specific to financial applications. It evaluates models’ understanding of financial discussions, extraction of key points, and effectiveness in condensing lengthy audio into concise summaries. This capability is crucial for stakeholders needing quick insights from extensive financial dialogues, such as investor briefings or regulatory meetings.
AudioLLMs can process audio with a maximum length of 30 seconds. Therefore, AudioLLM cannot perform the summarization task independently, we developed a processing pipeline. Given input data represented as $(A,Q,L)$, where AA denotes the long audio recording, $Q$ the prompt instruction, and $L$ the reference summary, the pipeline first segments the long audio into smaller chunks (see Section 2.1.2). Each chunk is transcribed by an AudioLLM into text segments $t_{i}$, which are combined into a complete transcription $T$. This transcription is then input into an LLM to generate the summary $S$. Summarization performance is evaluated using two widely adopted summarization evaluation metrics: Rouge-L [lin2004rouge] and BertScore [zhangbertscore], computed between the generated summary ($S$) and the reference summary ($L$). The higher values for these metrics indicate better AudioLLM performance.
2.2 FinAudio Datasets
The FINAUDIO benchmark contributes a unified framework by careful curation and task reformulation, transforming existing datasets into meaningful and comprehensive evaluation tasks. As summarized in Table 1, we construct MDRM-test and SPGISpeech-test by further developing their respective existing datasets for short ASR evaluation, create Earnings-21 and Earnings-22 test sets for long ASR tasks, and introduce our newly structured FinAudioSum for audio summarization task. Through this curation process, we establish over 430 hours of purposefully organized financial audio that addresses domain-specific evaluation needs not met by the original datasets. Detailed dataset descriptions are provided in Appendix B.
@@ 2.2 FinAudio Datasets
3 Experiment Results and Analysis
We describe the evaluated AudioLLMs and the experiment setup in the appendix C.
@@ 3.1 Results on ASR Tasks
Table 2 shows that Whisper-v3 achieves notably low WER scores (2%–3%) on both datasets, highlighting its robust speech recognition without domain-specific tuning. The Qwen2-Audio series (base and instruct versions), GTP-4o-audio, and Gemini exhibit moderate performance with WER scores between 4% and 6%. Conversely, SALMONN (7B and 13B) demonstrates substantially lower effectiveness, with WER around or exceeding 40%–50%. Despite SALMONN’s limited performance, these findings highlight the potential of other mainstream AudioLLMs to support voice-based financial services, particularly conversational assistants handling iterative short audio communications.
Datasets Models Whisper Qwen2-Audio Qwen2-Audio SALMONN SALMONN Gemini Gemini GPT-4o-audio -v3 -7B -7B-Instruct -7B -13B -1.5-flash -2.0-flash -transcribe MDRM-test 2.14 3.97 4.68 51.52 49.17 4.850 4.321 4.23 SPGISpeech-test 2.88 4.42 5.74 39.51 41.17 5.802 5.143 4.66 Earning-21 11.85 26.06 29.58 83.20 80.54 18.58 19.17 15.78 Earning-22 15.93 42.76 33.65 88.50 86.23 27.13 28.12 21.37
Table 2: Comparison of model performance based on WER across various speech datasets on ASR.
Results on long financial audio recording ASR.
@@
In the long financial audio ASR task, Whisper-v3 remained the top performance. It achieves lower WER in the range of approximately 12% - 16%. The GPT-4o-audio emerged as the second-best performer, with WER spanning roughly 15%–21%. The error rate of Qwen2-Audio series models has also increased significantly. SALMONN once again reported the highest error rates across both financial datasets (80–88%). These results reveal a decline in performance across all AudioLLMs, indicating that AudioLLMs face significant challenges when processing long financial audio data. The SALMONN series models performed poorly on both ASR tasks. Upon reviewing the experimental logs, we observed that SALMONN frequently produced empty outputs or failed to understand prompt instructions. This suggests that SALMONN might be overfitting to specific audio data or tasks, lacks robust instruction-following capabilities, and exhibits a limited understanding of financial audio content.
@@
In Figure 2, we find no significant difference in performance between Gemini-1.5 and Gemini-2.0. The open-source model Whisper-v3 achieves the best overall performance across all datasets. These results demonstrate robust capabilities even within long financial audio recognition. The Qwen-2 series, another open-source model, delivers comparable results to the closed-source Gemini models on short-audio ASR tasks but performs slightly worse on long-audio ASR. These findings demonstrate that open-source models can outperform their closed-source counterparts in financial audio scenarios, indicating that open-source AudioLLMs represent an effective, cost-efficient, and privacy-preserving solution for developing AI-driven financial products.
3.2 Results on summarization task and ablation study
For summarization, we evaluated Whisper-v3, Qwen2-Audio-7B-Instruct, Gemini-1.5-flash and GPT-4o-audio. The overall summarization quality depends on the ASR performance of AudioLLMs. Gemini-1.5-flash achieved the best ROUGE-L and BERTScore, while Whisper slightly outperformed Qwen2-Audio-7B-Instruct, showing strong general summarization ability without domain-specific tuning. We describe the detailed results in Appendix D
We also examined prompt robustness across ASR tasks. Whisper-v3 and Qwen2-Audio-7B-Instruct maintained stable performance under varied prompts, whereas Qwen2-Audio-7B was more sensitive, highlighting the importance of instruction tuning. Finally, our error analysis revealed that AudioLLMs often struggle with financial terminology and numerical information, underscoring the need for improved domain adaptation. We show the detailed analysis and visualization in Appendix E. We also conduct error analysis in Appendix F.
4 Conclusion and Research Outlook
We propose FinAudio, an AudioLLM Benchmark for financial audio data. We constructed a total of 400h+ of audio data for three tasks and evaluated it on seven AudioLLMs. Our systematic evaluation revealed substantial challenges, notably the pronounced difficulty models face in accurately processing long-form financial audio and specialized terminology. Future research directions for applying AudioLLMs in the financial industry include (1) extending the input context window length of AudioLLM, (2) enhancing the models’ comprehension of numerical transcription and specialized financial terminology, and (3) improving their reasoning capabilities in financial contexts.
Appendix A Related Work
A.1 Advancements in AudioLLM
Among the modalities addressed by multimodal LLMs, audio stands out as particularly critical and challenging [su2025audio]. Several multimodal models, such as Gemini Pro [team2023gemini] and PandaGPT [su2023pandagpt], have been adapted to process audio data. Prominent AudioLLMs include AudioGPT [huang2024audiogpt], SpeechGPT [zhang2023speechgpt], AudioPaLM [rubenstein2023audiopalm], Qwen-Audio [chu2023qwen], WavLLM [hu2024wavllm], Whisper [radford2023robust], Salmonn [tangsalmonn], and Qwen2-Audio [chu2024qwen2]. Correspondingly, general-domain benchmarks like AirBench [yang2024air], emphasizing conversational scenarios, and AudioBench [wang2024audiobench], evaluating eight tasks across twenty datasets, have been developed to assess AudioLLMs. However, these benchmarks contain limited financial audio data, inadequate for a comprehensive evaluation in financial contexts.
A.2 Progress in Multimodal Financial LLMs And Benchmark
Recent advances in general-domain LLMs have accelerated the emergence of specialized financial LLMs, such as FinGPT [liu2023fingpt], InvestLM [yang2023investlm], and BloombergGPT [wu2023bloomberggpt]. The progression toward multimodal financial LLMs—including Open-FinLLMs [xie2024open] and FinTral [bhatia2024fintral]—further highlights efforts to address complex financial tasks involving charts and tabular data. These multimodal models overcome the limitations of traditional approaches [ang2022guided, bhatia2024fintral, Yu2022DeepTF] by more flexibly adapting to financial contexts. As financial LLMs become increasingly specialized, the demand for robust evaluation grows. In response, various benchmarks have been proposed. FLUE [shah2022flue] introduced the first benchmark for financial NLP evaluation, covering tasks such as sentiment analysis, headline classification, named entity recognition, structure boundary detection, and question answering. PIXIU [xie2023pixiu] extended these evaluations into multimodal document analysis, while FinBen [xie2025finben] significantly expanded coverage, featuring 36 datasets across 24 financial tasks. Despite these advances, existing benchmarks remain predominantly text-focused, overlooking audio—an essential modality in financial applications.
Appendix B Dataset
B.1 Short financial audio clip datasets:
Two datasets provide financial audio clips: MDRM [qin2019you] and SPGISpeech [o2021spgispeech], both derived from earnings conference call recordings. The original audio data was segmented at the sentence level and aligned with corresponding transcripts. Each dataset is initially split into training and test subsets; we utilize only the test sets for evaluation.
  • MDRM [qin2019you]: This is the first dataset to use multimodal financial data for stock volatility and movement prediction. MDRM includes both audio recordings and text transcripts from the 2017 earnings calls of 500 major companies listed on the S&P 500 and traded on U.S. stock exchanges. There are a total of 280 companies with 572 earnings conference calls. The author further segmented earnings conference calls into individual sentences based on the text and conducted corresponding audio segments. There are a total of 88,829 audio clips. We split the dataset into training and test sets using a 75%–25% ratio, with the test portion utilized as the evaluation set for the benchmark. The MDRM-test set includes 22,208 audio clips totaling 87 hours.
  • SPGISpecch [o2021spgispeech]: The dataset comprises 5,000 hours of earnings conference call recordings. The original audio is segmented by sentence and professionally transcribed into a structured format. Each audio segment is meticulously segmented by sentences, ranging from 5 to 15 seconds in duration. Partitioned initially into training and test sets, our analysis focuses solely on the test subset, comprising 39,341 clips and approximately 130 hours of audio.
B.2 Long financial audio recording datasets:
There are two other publicly available earnings conference call datasets containing raw audio recordings, which can be leveraged for the long financial audio ASR task:
  • Earnings-21 [del2021earnings]: The dataset comprises 44 complete earnings conference call recordings from 2021 (approximately 39 hours). A distinguishing characteristic of Earnings-21 is its emphasis on named entities, financial jargon, and numerical data, which often pose difficulties for ASR models. The aligned transcripts are further enhanced with detailed formatting and entity tags, making it possible to conduct granular error analysis.
  • Earnings-22 [del2022earnings]: This is a updated version of Earnings-21. It has 125 audio samples with a total of 120 hours. Each audio recording includes a corresponding manual transcript. We use both datasets entirely for evaluation.
B.3 FinAudioSum
We introduce FinAudioSum, a novel dataset created explicitly for long-form financial audio summarization tasks, building upon the foundation of the ECTSum dataset [mukherjee2022ectsum]. While ECTSum was initially designed for summarizing earnings calls based solely on textual transcripts, our contribution uniquely expands this resource into the audio modality, addressing a critical gap in multimodal financial summarization research. ECTSum comprises 2,425 earnings transcripts paired with expert-generated, telegram-style summaries. We obtain corresponding audio recordings for the ECTSum test set from earningcast [^1] https://earningscast.com/. To ensure uniqueness and avoid redundancy, we carefully eliminate any overlapping audio entries from Earnings-21 and Earnings-22 datasets, covering the years 2019–2022. The final FinAudioSum dataset includes 64 recordings totaling 55 hours.
Task Input Instruction Output ASR for Short financial audio clip [Uncaptioned image] fast fashion has made it possible for everyone to have access to aesthetically thoughtful clothing (7s) Convert the audio speech into a text transcript. fast fashion has made it possible for everyone to have access to aesthetically thoughtful clothing ASR for Long financial audio recording [Uncaptioned image] Good morning, everyone, and welcome to the NextEra Energy Inc… (40min+) Transcribe the spoken words into written form. Good morning, everyone, and welcome to the NextEra Energy Inc… FinAudioSum [Uncaptioned image] Good morning, everyone, and welcome to the NextEra Energy Inc… (40min+) Listen to the audio and provide the text version. In past three month, NextEra Energy Inc release…
Table 3: Examples of data in the FinAudio evaluation suite.
Appendix C Model and Experiment Setup
We evaluate the performance of seven representative AudioLLMs on the FinAudio benchmark. The chosen models include five open-source models—Whisper [radford2023robust], Qwen2-Audio-7B, Qwen2-Audio-7B-instruct [chu2024qwen2], SALMONN-7B, and SALMONN-13B [tangsalmonn]—and three closed-source models, Gemini-1.5-flash and Gemini-2.0-flash [team2023gemini], and GPT-4o-audio. These models span various parameter sizes, open- and closed-source implementations, and include both base and instruction-tuned versions, enabling a comprehensive assessment of AudioLLMs for financial audio tasks. This comprehensive evaluation enables a thorough analysis of current AudioLLMs’ strengths and limitations within financial audio contexts.
To ensure fairness, we standardized the prompt as: “Convert the audio speech into a text transcript,”. For the summarization task, we employed GPT-4o (temperature=0) to generate summaries from the model-produced transcripts, ensuring a uniform summarization baseline. All open-source AudioLLM evaluations are conducted on two NVIDIA RTX A6000 GPUs, each equipped with 48GB of DRAM, necessitating approximately 200 hours per evaluation round. Additionally, to account for experimental variability and ensure the robustness of our assessments, each experiment was conducted three times, with the reported results representing the average performance.
Appendix D Results on summarization task